# Optimization of deterministic photonic graph state generation via local operations

### v2

@@ Acknowledgement
Figure S3: The circuit to generate the best alternative graph in the orbit of an RGS with $n=2m+8$ qubits. The elementary circuit block in the middle is repeated for $1\leq k\leq m$ while the rest of the circuit is fixed. For odd $m$: $V_{1}$ = X, $V_{2}$ = U, and $V_{3}$ = $U^{\prime}$, and for even $m$: $V_{1}$ = I, $V_{2}$ = $V_{3}$ = H. $U$ and $U^{\prime}$ are Clifford unitaries $HPHPX$ and $HPHP$ respectively.
B: Graph-to-circuit maps
@@
The optimization outcome is dependant on the specific deterministic graph to circuit map that is used to find a generating circuit for each input graph, however, the framework can take any such graph-to-circuit mapping protocol to work with. The first protocol we implemented (Ref. [27] in the main text) was selected to ensure that the algorithms utilized the minimum number of emitters necessary to generate the entanglement present in a graph state. However, the number of emitters may not always be the primary cost metric and thus other maps can become more suitable options depending on the platform and hardware. For instance, Kaur et al. [2] have shown that considering solely the photon loss as the dominant factor in the failure of an all-photonic repeater protocol, a map that allows for the use of extra emitters (over the theoretical minimum) actually results in a better overall success rate.
@@
Since the optimization scheme proposed here is protocol-agnostic, we analyzed repeater graph states of varying sizes using another specific algorithm (Algorithm 2, as referenced in [2]). In this analysis, we employed 4 emitters instead of the minimum 2. The results are promising, revealing patterns analogous to those observed previously. For example, when plotting the cnot-vs-edge count diagram (Fig. S4), two distinct parts of the orbit become evident, with one showing a strong linear correlation. It is worth noting that the strength of the employed algorithm is in its effectiveness in depth reduction, thus the number of cnots, even in the best case (scaling as $3\times\lfloor n/3\rfloor$ with $n$ being the size of the graph), is larger compared to our previously identified optimal case ($n/2-2$). However, when considering circuit depth (to be specific emitter depth as defined in the main text) the results show an advantage over the circuit depths obtainable with the previous protocol ($n/2+6$ vs $n$ scaling). For the 4-emitter scenario, the optimal graph featuring the least depth coincides with the case that required the least number of cnots. Additionally, the results can be straightforwardly generalize to arbitrary large RGS as we did before; no matter the size of the RGS, using orbit enumeration Algorithm (1), the optimal shape resulting in the above mentioned advantages in circuit depth, is always the $3^{rd}$ graph in the list of graphs added to the orbit in the algorithm.
@@
Figure S4: Number of cnots vs Number of edges (x-axis in logarithmic scale) correlation diagram of a 30-qubit RGS. Points correspond to graphs in the orbit. The orbit show two distinct behaviors, a group of members showing linear correlation while the other is constant in the number of cnots.
C: Repeater graph performance
@@
A repeater graph is made up of a set of core nodes forming a complete sub-graph, and a set of leaf nodes (arms) attached to each one of the core qubits. The final goal of a repeater state can be thought of as distilling a perfect maximally entangled EPR pair that is established between two distance parties by distributing the leaf qubits (Fig. S5). Therefore, the overall performance of a single repeater graph state can be determined by evaluating the quality of this distilled 2-qubit state, that can be quantified, for instance, via its fidelity with a maximally entangled state.
@@
Figure S5: EPR distillation process. From left to right, first a pair of qubits are selected to comprise the final 2-qubit state, the rest of the qubits, except the neighbors of the chosen pair are measured in the Z basis, and the two neighbors are measured in the X basis. The resultant state is locally equivalent to a graph where the two qubits are connected to each other, representing a maximally entangled state.
@@
We now estimate this final fidelity for two cases, a 10-qubit and a 20-qubit RGS, and for each case we consider 2 scenarios: (i) using the original generation circuit and (ii) using the optimized one with reduced cnots. Considering the emitter-emitter CNOT gates between to be dominant sources of imperfection, we applied a noise model with an added depolarizing channel (with a depolarizing rate of 1%) to each of the CNOT gates in the generation circuit. For each case, we simulate the distillation procedure for different choices of the two arms used to form the EPR pair, and report the lower-bound for the fidelity of the final state. The results are listed in the table Table S1, and indicate that the same fidelity for the non-optimized scenario can be achieved by using the optimized circuit for an RGS twice the size.
@@
Table S1: Fidelity of the final entangled pair distilled from a repeater graph state
@@
Simulating for a range of sizes and depolarization rates, the results remain similar, confirming that the optimization allows for doubling the size of the generated RGS state without affecting the performance of the repeater. Such increase in size can lead to orders of magnitude increase in the achievable quantum key distribution rate. Specifically, the repeater protocol’s success rate is given by $(1-(1-{P_{b}}^{n/4}))^{N}$ [3]. Here, $n$ is number of photons in RGS, $N$ is the number of repeater stations between the two parties, and $P$ is the probability of a successful Bell state measurement between two photons from adjacent repeater stations. Using the parameters from [3]’s realistic example cases, and considering a total distance of $400$ $km$, with $N=100$ and $P_{b}=0.416$, we find an $2.66\times 10^{5}$ increase in success rate when $n$ is doubled from $15$ to $30$ qubits.
The value for $P_{b}$ is obtained by multiplying the probability of both photons making it to the Bell state measurement (BSM) station by the success probability of the BSM itself ($=50\%$). Assuming a loss of $\eta=0.2$ $dB/km$ and $2$ $km$ travel distance for each photon from two adjacent RGS to reach the BSM station in between (total distance $=400/100=4$ $km$) the probability of having both photons is equal to $0.832$ and then $P_{b}=0.416$.
D: Correlation analysis
@@
In a correlation analysis of the local equivalency class of a graph, the goal is to determine whether there exists a significant correlation between a graph metric (a parameter related to the shape of a graph) and a cost parameter of its generating circuit. We examine how changes in the average values of one metric (chosen as the dependent variable) correspond to variations in the other metric (chosen as the independent variable) across the locally equivalent graphs in an orbit or in a sample from it. For instance, in Fig. S6, the independent variable (x-axis) the is set to be the number of edges in the graph and the dependent variable (y-axis) is chosen as the average of the circuit’s gate-count over the graphs sharing the same number of edges in the sample taken from the orbit. This diagram is derived for a sample of the LC orbit of a random graph of size $n$=20 with an edge probability of $p$=0.95. The Pearson correlation coefficient, used to measures the degree of linear relationship between two variables [4], for this specific diagram is $94\%$. Repeating the same procedure for 100 random graphs of the same type ($n$=20, $p$=0.95), we find the distribution of the correlation coefficients to be as seen in Fig. S7. This shows an average Pearson correlation of $0.86\pm 0.07$, indicating that for most of the graphs of this type, a strong correlation (over 0.5) holds true.
@@
Figure S6: The correlation diagram for the average number of unitary gates in a circuit vs the number of edges, over a sample of size 200 from the orbit of the graph. The error bars indicate one standard deviation when more than one graph is associated with the same number of edge in the sample.
@@
Figure S7: The distribution of Pearson coefficients for ‘edge count vs gate count’ correlations for 100 trials with different random graphs of size 20 and edge probability 0.95.
@@
With the correlation confirmed, one can then, in this case, aim for the graphs in the orbit with a smaller number of edges, corresponding to circuits with fewer total number of unitary gates. To do this, we developed an active edge reduction strategy to get to the sparser graphs in the orbit through applying local complementation operations on specific nodes, starting from the initial graph. In particular, we select the node with the highest product of node degree (number of neighbors) and clustering coefficient [5], and apply local complementation on it. The clustering coefficient of a node is defined as the ratio of the number of existing edges in the neighborhood of that node, to the maximum possible number of edges in that neighborhood. The process is iteratively repeated on the resulting graph until further steps would no longer reduce the number of edges.
@@
This active edge reduction strategy shows a one to two order of magnitude optimization speed up compared with a random search method, where a random sample from the graphs in the orbit is taken, converted into corresponding circuits, and the best circuit (with minimum unitary gate count) is selected. Fig. S8 shows a runtime scaling diagram for graphs of size n=30 to 55 while the sample size is chosen so that the performance of the both methods is similar as seen in Fig. S9.
@@
The edge reduction strategy used here take at most $\mathcal{O}(n^{2})$ steps, as in each step, we require for the number of edges to be reduced at least by one, and there are at most $n(n-1)/2$ edges in a graph. The clustering coefficient of the graph needs to be calculated in each step, which has a time complexity of $\mathcal{O}(n^{3})$ for a brute-force algorithm. Therefore, the total complexity will be upper bounded by $\mathcal{O}(n^{5})$. However, in practice, the actual scaling with size is much smaller than the worst case scenario, for example, in the specific use case of dense graphs (edge probability 95%), the scaling turns out to be $n^{2.68}$ for our python implementation (see Fig. S10). The method may thus be used to for optimization of graphs with thousands of qubits within minutes on commercial computers.
@@
Figure S8: Runtime comparison for the same performance between the active edge reduction method (red) and a random search method (green). Error bars are 1 standard deviation for data averaged over 100 trials.
@@
Figure S9: Unitary gate count for optimization with the active edge reduction strategy (red), the random search method (green), compared to the initial gate count(black). Error bars are 1 standard deviation for data averaged over 100 trials.
@@
Figure S10: Polynomial scaling of the runtime with graph size for the proposed edge reduction strategy. Error bars are 1 standard deviation for data averaged over 100 trials.
E: Exhaustive analysis of small graphs
@@
To showcase the results of the exhaustive analysis, we look into an entanglement class associated with the graph (A) in Fig. S11, i.e., all graph states that are LC equivalent to this graph up to relabeling of the nodes. We have identified multiple intermediate target graphs to utilize in the generation process of any member of this entanglement class, each tailored for optimizing a certain cost parameter. In Fig. S11, examples of a case with the least number of cnots (B), a case with the smallest emitter depth (C), and a case with minimal number of unitary gates (D) are presented. More details for each option are listed in Table S2. We remark that there might be more than one best performing intermediate graphs depending on the cost function used. One can find the maximum improvement for each cost parameter by comparing the best values with the worst ones present in the last column, showing the maximum of the cost parameters over this class when no optimization is done. Overall, any arbitrary cost function can be employed to filter the results of such an exhaustive analysis, e.g, the fidelity of the final state with the target when a proper noise model is employed. The full dataset containing the results from the analysis of all 45 entanglement classes of graph with $n\leq 7$ nodes is available at Ref. [6].
@@
Figure S11: A set of optimal intermediate graphs (B-D) for generation of the entanglement class associated with graph (A), minimizing different cost functions, in the same order from (B) to (D): cnot count, emitter depth, and unitary gate count. There might be more than one best graph for each metric in the orbit. Node labels indicate the photon emission order.
Table S2: Summary of the cost metrics for the alternative graphs outlined in Fig. S11.
@@
Depth [^1] Emitter depth as defined in the main text 14 23 7 10 41

### v3

@@ IV.1 Repeater graph states
Figure 6: The circuit to generate the best alternative graph in the orbit of an RGS with $n=2m+8$ qubits. The elementary circuit block in the middle is repeated for $k$ time with $1\leq k\leq m$ while the rest of the circuit is fixed. $U$ and $U^{\prime}$ are Clifford unitaries $HPHPX$ and $HPHP$ respectively. For odd $m$: $V_{1}$ = X, $V_{2}$ = U, and $V_{3}$ = $U^{\prime}$, and for even $m$: $V_{1}$ = I, $V_{2}$ = $V_{3}$ = H.
Such generalization allows us to navigate the orbit to reach the graph with the minimum cnot count. This also provides, for the first time, an efficient polynomial-time algorithm to generate and access the entire orbit of an arbitrarily large RGS. The procedure is outlined in detail in Fig. 5 below. It is noteworthy that the LC orbit exploration via brute-force search methods is not an efficient and scalable process in general, as one needs to perform a search over all possible sequences of local complementation operations on graph nodes and repeat the whole process for every new graph found in each round. In fact, even determining the size of the orbit for an arbitrary graph is a #P-complete problem [39].
By looking into the generating circuit for each of the graphs in the orbit, one can find different elementary circuit blocks that are to be utilized to add photons to the state in a sequential manner, constituting different parts of the target structure. The generation circuit blocks in Fig. 3 (d-e) demonstrate how the discussed graph transformations can change the scaling of the required number of 2-qubit gates from $n$ to $n/2$ for an $n$-qubit RGS, since the emission of each photon pair belonging to the transformed part of the graph requires a single cnot gate as opposed to the two gates needed when using the original RGS shape. Ultimately, the best (with respect to the use of cnot gates) alternative graph in the orbit can be identified by continuing this observed trend, and is achieved by sequentially applying the LC rule to all central nodes of an RGS, i.e., LC rule on nodes $\{1,3,\ldots,n-1\}$, if the number of arms ($n/2$) in the RGS is an odd number. For even $n/2$, the last node in the sequence is replaced by the leaf node connected to it, resulting in the set $\{1,3,\ldots,n-2\}$. The cnot count is, as a result, reduced from $n-3$ to $n/2-2$. The optimal cases are depicted in Fig. 5 and their corresponding circuits are shown in Fig. 6. This optimized scaling, identified systematically here, aligns with the scaling of the ad-hoc protocol described in Ref. [24], which was specifically designed for the optimal generation of RGS using a different approach and with another emission sequence. Note that the identified pattern and the optimization result are subject to change when different graph-to-circuit mapping protocol are employed. In Appendix A, we expand on how using another graph state generation protocol outlined in Ref. [28] results in scaling advantages in terms of circuit depth for RGS generation.
An important benefit of reducing the number of cnot gates, aside from mitigating the technical complications of realizing the generation circuit, is that it nearly doubles the maximum size of the producible RGS. This is because the size of the state is primarily limited by the emitter’s coherence time and how many emission events can be performed before the emitter qubits lose coherence, and the main time consuming operations in the circuit are the emitter-emitter cnots [28]. The significance of this size increase is better reflected in the repeater protocol’s success probability in establishing an EPR pair between two distant parties using multiple repeater stations in between, which is given by $P=(1-(1-P_{b})^{n/4})^{N}$ [9]. Here, $n$ is number of photons in the used RGS, $N$ is the number of RGS stations between the two parties, and $P_{b}$ is the probability of performing a successful entanglement swapping between two photons coming from two adjacent repeater stations. Doubling the size from $n$ to $2n$, especially in the realistic long distance regime with $n$ being in the order of tens and $N$ in hundreds, results in several orders-of-magnitude increase in the overall success rate. See Appendix B for more details on how the optimization affects the size and performance of the RGS.
IV.2 Non-structured graphs
To demonstrate the applicability of our method beyond structured graphs, we also investigated the opposite end of the spectrum, i.e., random graphs [40]. In order to expect a generalizable optimization behavior, a shared trait is needed in the set of graphs under analysis. To this end, we limit the study to random dense cases, specifically focusing on graphs with a 95% edge probability between each pair of nodes. Conjecturing a correlation between the generation cost and the shape of a graph, we chose a simple metric to parameterize the graph shape: the number of edges. The existence of such a correlation can be directly confirmed for graphs of small size, e.g., with 15 nodes, as the whole orbit is accessible by exhaustive search. By choosing the graph with the least edge count in the orbit, we observed average reductions of $65\%\pm 11\%$ in the total number of unitary gates of the circuit and $46\%\pm 18\%$ in emitter depth. One can efficiently test this correlation for larger graphs, as there are efficient methods to find sparse graphs within the orbit of dense graphs; see Appendix C for one such efficient strategy developed by the authors, with no claim of being optimal. Figure 7 shows the cost reduction achieved when extending the use of this strategy for graph of up to 65 nodes.
We also conducted an exhaustive analysis (see Appendix D) of graphs with up to 7 qubits [41], enumerating all the possible 45 entanglement classes as defined in Ref. [33] and comparing the generation circuits across $\sim 1.9\times 10^{6}$ labeled connected graphs [42].
All the presented analyses were conducted on a 3.2 GHz, 8-core, Apple M1 Pro chip with 16 GB of RAM, using our Python package “GraphiQ” [43].
V Discussion
We introduce an optimization framework tailored for the deterministic generation of photonic graph states, a task that poses significant theoretical and experimental challenges due to its inherent constraints. Our approach leverages the concept of local Clifford equivalency and local complementation to effectively identify alternative graphs in the same local equivalency class that can serve as more resource-efficient paths for reaching a target state.
It is worth noting that the employed cost metrics can be tailored to meet the requirements of specific experimental platforms, e.g., as we have done with the ‘emitter depth’ to address the limitations in emitters’ coherence time. Additionally, the application can be easily extended to the probabilistic fusion-based graph generation protocols, such as the one in Ref. [5], by facilitating the expansion of both the size and diversity of resource states that are required as building blocks in making larger graphs.
Acknowledgement
Appendix A: Graph-to-circuit maps
@@ Acknowledgement
In general, the optimization outcome is dependent on the specific deterministic graph to circuit map that is used to find a generating circuit for each input graph. However, the framework can take any such graph-to-circuit mapping protocol to work with. The first protocol we implemented (Ref. [27]) was selected to ensure that the algorithms utilized the minimum number of emitters necessary to generate the entanglement present in a graph state. However, the number of emitters may not always be the primary cost metric and thus other maps can become more suitable options depending on the platform and hardware. For instance, Kaur et al. [28] have shown that considering solely the photon loss as the dominant factor in the failure of an all-photonic repeater protocol, a map that allows for the use of extra emitters (over the theoretical minimum) actually results in a better overall success rate.
@@
Since the optimization framework proposed here is applicable to other protocols, we analyzed repeater graph states of varying sizes using another specific algorithm (Algorithm 2, as referenced in [28]). In this analysis, we employed four emitters instead of the minimum two required for RGS generation. The results are promising, revealing patterns analogous to those observed previously. For example, when plotting the cnot-vs-edge count diagram (Fig. 8), two distinct parts of the orbit become recognizable, with one showing a strong linear correlation. It is worth noting that the strength of the employed algorithm is in its effectiveness in depth reduction, thus the number of cnots, even in the best case (scaling as $3\times\lfloor n/3\rfloor$ with $n$ being the size of the graph), is larger compared to our previously identified optimal case ($n/2-2$). However, when considering circuit depth (to be specific emitter depth as defined in the main text) the results show an advantage over the circuit depths obtainable with the previous protocol ($n/2+6$ vs $n$ scaling). For the four-emitter scenario, the optimal graph featuring the least depth coincides with the case that requires the fewest cnot gates. Additionally, the results can be straightforwardly generalized to arbitrary large RGS as we did before; no matter the size of the RGS, when using the orbit enumeration Algorithm (5), the optimal shape resulting in the above-mentioned advantages is always the third graph in the list of graphs added to the orbit via the algorithm.
@@
Figure 8: Number of cnots vs Number of edges (x-axis in logarithmic scale) correlation diagram of a 30-qubit RGS. Points correspond to graphs in the orbit. The orbit shows two distinct behaviors, a group of members showing a linear correlation while the others are constant in the number of cnots.
Appendix B: Repeater graph performance
@@
A repeater graph is made up of a set of core nodes forming a complete sub-graph, and a set of leaf nodes (arms) attached to each one of the core qubits. The final goal of a repeater state can be thought of as distilling a perfect maximally entangled EPR pair that is established between two distant parties by distributing the leaf qubits (Fig. 9). Therefore, the overall performance of a single repeater graph state can be determined by evaluating the quality of this distilled 2-qubit state that can be quantified, for instance, via its fidelity with a maximally entangled state.
@@
Figure 9: EPR distillation process. From left to right, first a pair of qubits is selected to comprise the final two-qubit state, the rest of the qubits, except for the two nodes connected to the chosen pair are measured in the $Z$ basis, and finally the two neighboring nodes are measured in the $X$ basis. The resultant state is locally equivalent to a graph where the two qubits are connected to each other directly, representing a maximally entangled state.
@@
We now estimate this final fidelity for two cases, a 10-qubit and a 20-qubit RGS, and for each case we consider 2 scenarios: (i) using the original generation circuit and (ii) using the optimized one with reduced cnots. Considering the emitter-emitter cnot gates to be the dominant source of imperfection, we applied a noise model with an added depolarizing channel (with a depolarization rate of 1%) to each of the cnot gates in the generation circuit. For each case, we simulate the distillation procedure for different choices of the two arms used to form the EPR pair, and report the lower-bound for the fidelity of the final state. The results are listed in the table Table 1 and indicate that the same fidelity for the non-optimized scenario can be achieved by using the optimized circuit for an RGS twice the size.
@@
Table 1: Fidelity of the final entangled pair distilled from a repeater graph state
@@
Simulating for a range of sizes and depolarization rates, the results remain consistent, confirming that the optimization allows for doubling the size of the generated RGS state without affecting the performance of the repeater.
@@
Such an increase in size can lead to orders of magnitude increase in the achievable quantum key distribution rate. Specifically, the repeater protocol’s success rate is given by $(1-(1-{P_{b}}^{n/4}))^{N}$ [9]. Here, $n$ is the number of photons in the RGS, $N$ is the number of repeater stations between the two parties, and $P$ is the probability of a successful Bell state measurement between two photons from adjacent repeater stations. Using the parameters from [9]’s realistic example cases, and considering a total distance of $400$ $km$, with $N=100$ and $P_{b}=0.416$, we find a $2.66\times 10^{5}$ increase in success rate when $n$ is doubled from $15$ to $30$ qubits. The value for $P_{b}$ was obtained by multiplying the probability of both photons making it to the Bell state measurement (BSM) station, by the success probability of the BSM itself ($=50\%$ [12]). Assuming a loss of $0.2$ $dB/km$ and $2$ $km$ travel distance for each photon from two adjacent RGS to reach the BSM station in the middle (half the total distance $=400/100=4$ $km$ between two repeaters) the probability of having both photons present is equal to $0.832$ and so $P_{b}=0.5\times 0.832=0.416$.
Appendix C: Correlation analysis
@@
In a correlation analysis of the local equivalency class of a graph, the goal is to determine whether there exists a significant correlation between a graph metric (a parameter related to the shape of a graph) and a cost parameter of its generating circuit. We examine how changes in the average values of one metric (chosen as the dependent variable) correspond to variations in the other metric (chosen as the independent variable) across the locally equivalent graphs in an orbit or in a sample from it. For instance, in Fig. 10, the independent variable (x-axis) is set to be the number of edges in the graph and the dependent variable (y-axis) is chosen as the average of the circuit’s gate-count over the graphs sharing the same number of edges in the sample taken from the orbit. This diagram is derived for a sample of the LC orbit of a random graph of size $n$=20 with an edge probability of $p$=0.95. The Pearson correlation coefficient, used to measure the degree of linear relationship between two variables [44], for this specific diagram is $94\%$. Repeating the same procedure for 100 random graphs of the same type ($n$=20, $p$=0.95), we find the distribution of the correlation coefficients, as seen in Fig. 11. This shows an average Pearson correlation coefficient of $0.86\pm 0.07$, indicating that for most of the graphs of this type, a strong correlation—Pearson coefficient over 0.5—holds true.
@@
Figure 10: The correlation diagram for the average number of unitary gates in a circuit vs the number of edges, over a sample of size 200 from the orbit of the graph. The error bars indicate one standard deviation when more than one graph is associated with the same number of edges in the sample.
@@
Figure 11: The distribution of Pearson coefficients for ‘edge count vs gate count’ correlations for 100 trials with different random graphs of size 20 and edge probability 0.95.
@@
With the correlation confirmed, one can, in this case, aim for the graphs in the orbit that have a smaller number of edges, corresponding to circuits with fewer unitary gates overall. To do this, we developed an active edge reduction strategy to get to the sparser graphs in the orbit through applying local complementation operations on specific nodes, starting from the initial graph. In particular, we select the node with the highest product of node degree (number of neighbors) and clustering coefficient [45], and apply local complementation on it. The clustering coefficient of a node is defined as the ratio of the number of existing edges in the neighborhood of that node, to the maximum possible number of edges in that neighborhood. The process is iteratively repeated on the resulting graph until further steps would no longer reduce the number of edges.
@@
This active edge reduction strategy shows a one to two order of magnitude optimization speed up compared with a random search method, where a random sample from the graphs in the orbit is taken, converted into corresponding circuits, and the best circuit (with minimum unitary gate count) is selected. Fig. 12 shows a runtime scaling diagram for graphs of size n=30 to 55 while the sample size is chosen so that the performance of both methods is similar for each graph size. This can be seen in Fig. 13.
@@
The edge reduction strategy used here takes at most $\mathcal{O}(n^{2})$ steps to complete. This is because in each step we require the number of edges to be reduced at least by one, and there are at most $n(n-1)/2$ edges in a graph. The clustering coefficient of the graph needs to be calculated in each step, which has a time complexity of $\mathcal{O}(n^{3})$ using a brute-force algorithm. Therefore, the total complexity will be upper-bounded by $\mathcal{O}(n^{5})$. However, in practice, the actual scaling with size is much smaller than the worst case scenario, for example, in the specific use case of dense graphs (edge probability 95%), the scaling turns out to be $n^{2.68}$ for our python implementation (see Fig. 14). The method may thus be used for optimization of graphs with thousands of nodes within minutes on commercial processors.
@@
Figure 12: Runtime comparison for the same performance between the active edge reduction method and a random search method. Error bars are 1 standard deviation for data averaged over 100 trials.
@@
Figure 13: Unitary gate count for optimization with the active edge reduction strategy (red diamond), and the random search method (green circle). The sample size for random search is chosen such that the performance would be similar to the active method. The initial non-optimized gate count (black square) is also depicted for comparison. Error bars represent one standard deviation for each data point averaged over 100 trials.
@@
Figure 14: Polynomial scaling of the runtime with graph size for the proposed edge reduction strategy is evident from the line of best fit. Error bars represent one standard deviation for each data point averaged over 100 trials.
Appendix D: Exhaustive analysis of small graphs
@@
To showcase the results of the exhaustive analysis, we look into an entanglement class associated with the graph (A) in Fig. 15, i.e., all graph states that are LC equivalent to this graph up to relabeling of the nodes. We have identified multiple intermediate target graphs to utilize in the generation process of any member of this entanglement class, each tailored for optimizing a certain cost parameter. In Fig. 15, examples of a case with the fewest number of cnots (B), a case with the smallest emitter depth (C), and a case with minimal number of unitary gates (D) are presented. More details for each option are listed in Table 2. We remark that there might be more than one best performing intermediate graphs depending on the cost function used. One can find the maximum improvement for each cost parameter by comparing the best values with the worst ones present in the last column, showing the maximum of the cost parameters over this class when no optimization is done. Overall, any arbitrary cost function can be employed to filter the results of such an exhaustive analysis, e.g., the fidelity of the final state with the target when a proper noise model is employed. The full dataset containing the results from the analysis of all 45 entanglement classes of graphs with $n\leq 7$ nodes is available at Ref. [41].
@@
Figure 15: A set of optimal intermediate graphs (B-D) for the generation of the entanglement class associated with graph (A), minimizing different cost functions in the following order from (B) to (D): cnot count, emitter depth, and unitary gate count. There might be more than one best graph for each metric in the orbit. Node labels indicate the photon emission order.
Table 2: Summary of the cost metrics for the alternative graphs outlined in Fig. 15.
@@
Depth [^1] Emitter depth as defined in Section II 14 23 7 10 41